OpenStudy (amistre64):

what are the next 2 terms in this sequence? 6 9 14 22 36 57 88 __ __

OpenStudy (anonymous):

88 + 42 = 130

OpenStudy (amistre64):

88 is right, I had to dbl chk meself sothat there was no typos and I could gradeja lol

OpenStudy (anonymous):

130 + 56 = 186

OpenStudy (anonymous):

Am I right? :)

OpenStudy (amistre64):

Ana, very good :)

OpenStudy (anonymous):

^^

OpenStudy (anonymous):

I'm missing the sequence...

OpenStudy (amistre64):

teh key is in taking the differences down until they are the same ... takes about 3 or 4 cycles

OpenStudy (anonymous):

I think my brain exploded after reteaching myself (or learning for the first time...don't remember) Lagrange Multipliers

OpenStudy (anonymous):

The differences between one term to the other are: 6 9 14 22 36 57 88 __ __ 3 5 8 14 21 31 And the differences between one of each of these terms to the others are: 2 3 6 7 10 And finally: 1 3 1 3 So to get the next number you'll add, you add 1 to 10, 11. Then 11 to 31, 42, then 42 to 88 = 130. For the next, 3 + 11 = 14, 14 + 42 = 56, 130 +56 = 186.

OpenStudy (amistre64):

i started with 1 3 1 3 ; and added 2 to the next step to get: 2, 2+1 = 3, 3+3=6, 6+1=7, 7+3=10 ; then i picked a random number to start the next line; 3 .... yeps :)

OpenStudy (anonymous):

You hear that? That was the sound of my last 2 brain cells dying away :)

OpenStudy (amistre64):

theres also a fan method; where you go so far and then take the differences from the left side ... but i cant recall it to well

OpenStudy (anonymous):

amistre...if you get a chance, can you check #/groups/mathematics/updates/4dff6b430b8bbe4f12e7791e to make sure I got it right?